### v1

@@ 2.1 Probabilistic Structures
    ${\mathscr{N}}=(N^{0},N)$, where $N^{0}$ and $N$ are Poisson random measures associated to stationary $\mathbb{F}$-Poisson stationary point processes with characteristic jump measures $n(\mathrm{d}\zeta)$ and $n^{0}(\mathrm{d}\zeta)$, respectively.
We assume that the corresponding point processes $p^{0}$ and $p$ have $\mathbb{P}$-almost surely disjoint domains, or equivalently their jump times are almost surely disjoint. Finally, let ${\mathscr{G}}$ denote the completion of the sigma-algebra generated by the common noise processes $(W^{0},N^{0})$.
@@ 2.2 Controlled State Process
$(b,\sigma,\sigma_{0}):[0,T]\times{\mathscr{D}}_{n}\times U\mapsto\mathbb{R}^{n}\times{\mathscr{M}}_{n\times d}\times{\mathscr{M}}_{n\times\ell},$
$(\lambda,\lambda^{0}):[0,T]\times{\mathscr{D}}_{n}\times U\times E\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n},$
@@
All coefficients are measurable and there exist constants $K,L>0$ and $K(\zeta),L(\zeta)$ such that for any $(t,x,a,\zeta)\in[0,T]\times{\mathscr{D}}_{n}\times U\times E$ we have,
@@
and the constants $L(\zeta),K(\zeta)$ satisfy $\int_{E}(L(\zeta)^{p}+K(\zeta)^{p})\,(n^{0}+n)(\mathrm{d}\zeta)<\infty$ for any $p\geq 2$.
@@
$X^{\alpha}_{t}=X_{0}+\int_{0}^{t}$ $b(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}s+\int_{0}^{t}\boldsymbol{\sigma}(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}{\mathscr{W}}_{s}$
$+\int_{0}^{t}\int_{E}\boldsymbol{\lambda}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\widetilde{\mathscr{N}}(\mathrm{d}s,\mathrm{d}\zeta),\quad t\geq 0.$ (SDE$(X_{0},\alpha)$)
Under these assumptions, the standard Picard iteration technique implies the following theorem, see [42, Theorem 3.1].
@@
Under Assumption 2.3, for any probability basis $\gamma$, admissible control $\alpha\in{\mathscr{A}}(\gamma)$, and $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, there exists a unique $\mathbb{F}$-adapted càdlàg process $X^{\alpha}$ that satisfies SDE$(X_{0},\alpha)$. Furthermore, $\mathbb{E}^{\mathbb{P}}[\sup_{0\leq s\leq t}|X_{s}^{\alpha}|^{2}]<\infty$ for all $t\geq 0$.
@@ 3 Main Result
3.1 The Mean-Field Control Problem
Let a probability basis $\gamma$ be given and fix $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$. For a finite horizon $T<\infty$, let
$F:[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\mapsto\mathbb{R},\qquad G:{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\mapsto\mathbb{R},$
be given running and terminal cost functions satisfying Assumption 3.5 below. Let $J(X_{0},\alpha)$ be as in (1.3), and recall that ${\mathscr{G}}$ is the completion of the $\sigma$-algebra generated by the common noise processes $(W^{0},N^{0})$. In the absence of common noise, ${\mathscr{G}}$ is the trivial $\sigma$-algebra. We have the following important technical observation.
@@ 3.1 The Mean-Field Control Problem
The ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ measure-valued process
${\mathcal{L}}^{\alpha}_{t}={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,{\mathscr{G}})={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,\sigma(W^{0}_{s},N^{0}_{s}([0,s]\times B):s\leq t,B\in{\mathscr{E}}))$
has an $\mathbb{F}$-progressively measurable modification, and we always consider this version.
@@
Equipped with the product of the Skorokhod $J_{1}$ and Euclidean topologies, ${\mathscr{D}}_{n}\times U$ is a separable metric space. Since ${\mathscr{F}}_{0}$ is assumed to contain all $({\mathscr{F}},\mathbb{P})$-null sets, ${\mathcal{L}}^{\alpha}_{t}$ is adapted to the filtration $\mathbb{F}$. Therefore, there exists a $\mathbb{F}$-progressively measurable modification of ${\mathcal{L}}^{\alpha}$. ∎
For a given $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, the value function of the MFC is defined by,
$v(X_{0},\gamma):=\inf_{\alpha\in{\mathscr{A}}}J^{\gamma}(\alpha,X_{0})=\inf_{\alpha\in{\mathscr{A}}(\gamma)}\mathbb{E}^{\gamma}\left[\int_{0}^{T}F(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},{\mathcal{L}}^{\alpha}_{t})\,\mathrm{d}t+G(X^{\alpha}_{\cdot\land T},{\mathcal{L}}^{\alpha}_{T})\right].$ (3.1)
Further, for $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, let $\Gamma(\varrho)$ be the set of all bases $\gamma$ such that there exists $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0}^{\gamma})$ with distribution $\varrho$, let ${\mathcal{I}}^{\gamma}(\varrho)$ be the set of all such initial conditions, and we set
$V(\varrho):=\inf_{\gamma\in\Gamma(\varrho)}\inf_{X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)}v(X_{0},\gamma)=\inf_{\gamma\in\Gamma(\varrho)}\inf_{X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)}\inf_{\alpha\in{\mathscr{A}}}J^{\gamma}(\alpha,X_{0}).$ (3.2)
@@
We say that an MFC problem is invariant under the probability bases, if for any $\gamma,\tilde{\gamma}$, $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, $\tilde{X}_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\tilde{\gamma}}_{0})$, and $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ we have,
${\mathcal{L}}^{\gamma}(X_{0})={\mathcal{L}}^{\tilde{\gamma}}(\tilde{X}_{0})=\varrho\quad\Rightarrow\quad v(X_{0},\gamma)=v(\tilde{X}_{0},\tilde{\gamma})=V(\varrho).$
In the diffusion setting, invariance is proved in [20, Theorem 3.6] for ‘rich enough’ probabilistic structures, see [20, Section 2.1.3], and in [26, Proposition 2.4], and our Theorem 7.2 extends it to more general models. Although almost all problems have this property, some natural continuity and growth assumptions are needed as summarized in Assumption 7.1 and the counter example provided in [20, Section 3] highlights their necessity.
@@
3.2 The Potential Mean-Field Game of Controls
In this subsection, we give the definition of the potential mean-field game of controls equilibrium under common noise; see [13, Definition 2.1] for MFGs under common noise and [23, 24] for general MFC and MFG of controls close to our setting.
In these problems, the representative player takes an $\mathbb{F}$-adapted measure valued process $t\mapsto\nu^{*}_{t}\in{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ as given, and solves the optimal control problem of minimizing,
$J_{g}(X_{0},\alpha,\nu^{*}):=\mathbb{E}\left[\int_{0}^{T}f(t,X^{\alpha}_{\cdot\land t},\alpha_{t},\nu^{*}_{t})\,\mathrm{d}t+g(X^{\alpha}_{\cdot\land T},\nu^{*}_{T})\right]$ (3.3)
over all $\alpha\in{\mathscr{A}}$, where $f$, $g$ are as in (1.4). Then, the Nash equilibrium is defined as follows.
@@ 3.2 The Potential Mean-Field Game of Controls
    $\nu^{*}_{t}={\mathcal{L}}^{\alpha^{*}}_{t}={\mathcal{L}}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,{\mathscr{G}}),\qquad\mathbb{P}{\text{-a.s.}}\ \forall\ t\in[0,T].$ (3.4)
@@
In the absence of common noise, $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ is a deterministic ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$-valued process. Then, the second condition above reduces to $\nu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})$ for all $t\in[0,T]$. In this case, Borel measurability of $t\mapsto{\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})$ follows directly from the measurability of $(X^{\alpha^{*}},\alpha^{*})$.
@@ 3.3 Main Result
We assume that $F$,$G$ are measurable and they satisfy:
  • (Differentiability). The functions $F(t,x,a,\cdot)$ and $G(x,\cdot)$ are linearly differentiable.
  • (Quadratic Growth). For all $(t,x,x^{\prime},\nu,a)\in[0,T]\times{\mathscr{D}}_{n}^{2}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\times U$,
    $|F(t,x,a,\nu)|+|G(x,\nu)|+$ $\left|\delta_{\nu}F(t,x,a,\nu)(x^{\prime})\right|+\left|\delta_{\nu}G(x,\nu)(x^{\prime})\right|$
    $\leq C(1+\|x\|_{T}^{2}+|a|^{2}+\mathfrak{s}(\nu)+\|x^{\prime}\|_{T}^{2}).$
We prove the main result under the assumption of invariance proved in Theorem 7.2 under Assumption 7.1. Additionally, its proof reveals that it holds under a weaker version of invariance as discussed in Remark 5.3 below.

### v2

@@ 2.1 Probabilistic Structures
    ${\mathscr{N}}=(N^{0},N)$, where $N^{0}$ and $N$ are Poisson random measures associated to stationary $\mathbb{F}$-Poisson stationary point processes with characteristic jump measures $n^{0}(\mathrm{d}\zeta)$ and $n(\mathrm{d}\zeta)$, respectively.
We assume that the corresponding point processes $p^{0}$ and $p$ have $\mathbb{P}$-almost surely disjoint domains, or equivalently their jump times are almost surely disjoint. We further assume that $({\mathscr{W}},{\mathscr{N}})$ are independent. Finally, let ${\mathscr{G}}$ denote the completion of the sigma-algebra generated by the common noise processes $(W^{0},N^{0})$.
@@ 2.2 Controlled State Process
$(b,\sigma,\sigma_{0}):[0,T]\times{\mathscr{D}}_{n}\times U\to\mathbb{R}^{n}\times{\mathscr{M}}_{n\times d}\times{\mathscr{M}}_{n\times\ell},$
$(\lambda,\lambda^{0}):[0,T]\times{\mathscr{D}}_{n}\times U\times E\to\mathbb{R}^{n}\times\mathbb{R}^{n},$
@@
All coefficients are measurable and there exist constants $K,L>0$ and $K(\zeta),L(\zeta)$ such that for any $(t,x,x^{\prime},a,\zeta)\in[0,T]\times{\mathscr{D}}_{n}^{2}\times U\times E$ we have,
@@
and the constants $L(\zeta),K(\zeta)$ satisfy $\int_{E}(L(\zeta)^{2}+K(\zeta)^{2})\,(n^{0}+n)(\mathrm{d}\zeta)<\infty$.
@@
$\begin{split}X^{\alpha}_{t}=X_{0}+\int_{0}^{t}&b(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}s+\int_{0}^{t}\boldsymbol{\sigma}(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}{\mathscr{W}}_{s}\\ &+\int_{0}^{t}\int_{E}\boldsymbol{\lambda}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\widetilde{\mathscr{N}}(\mathrm{d}s,\mathrm{d}\zeta),\quad t\geq 0.\end{split}$ (SDE$(X_{0},\alpha)$)
Under these assumptions, the standard Picard iteration technique implies the following theorem, see [45, Theorem 3.1].
@@
Under Assumption 2.3, for any probability basis $\gamma$, admissible control $\alpha\in{\mathscr{A}}(\gamma)$, and $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, there exists a unique $\mathbb{F}^{\gamma}$-adapted càdlàg process $X^{\alpha}$ that satisfies SDE$(X_{0},\alpha)$. Furthermore, $\mathbb{E}^{\gamma}[\sup_{0\leq s\leq t}|X_{s}^{\alpha}|^{2}]<\infty$ for all $t\geq 0$.
@@ 3 Main Result
3.1 The Mean Field Control Problem
Let a probability basis $\gamma$ be given and fix $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0}^{\gamma})$. For a finite horizon $T<\infty$, let
$F:[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\to\mathbb{R},\qquad G:{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\to\mathbb{R},$
be given running and terminal cost functions satisfying Assumption 3.5 below. Let $J(X_{0},\alpha)$ be as in (1.2), and recall that ${\mathscr{G}}^{\gamma}$ is the completion of the $\sigma$-algebra generated by the common noise processes $(W^{\gamma,0},N^{\gamma,0})$. In the absence of common noise, ${\mathscr{G}}^{\gamma}$ is the trivial $\sigma$-algebra. We have the following important technical observation.
@@ 3.1 The Mean Field Control Problem
For any $\alpha\in{\mathscr{A}}(\gamma)$ the ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ measure-valued process
${\mathcal{L}}^{\gamma,\alpha}_{t}={\mathcal{L}}((X^{\gamma,\alpha}_{\cdot\land t},\alpha_{t})\,|\,{\mathscr{G}}^{\gamma})={\mathcal{L}}((X^{\gamma,\alpha}_{\cdot\land t},\alpha_{t})\,|\,\sigma(W^{\gamma,0}_{s},N^{\gamma,0}_{s}([0,s]\times B):s\leq t,B\in{\mathscr{E}}))$
has an $\mathbb{F}^{\gamma}$-progressively measurable modification, and we always consider this version.
@@
Equipped with the product of the Skorokhod $J_{1}$ and Euclidean topologies, ${\mathscr{D}}_{n}\times U$ is a separable metric space. Since ${\mathscr{F}}^{\gamma}_{0}$ is assumed to contain all $({\mathscr{F}}^{\gamma},\mathbb{P}^{\gamma})$-null sets, ${\mathcal{L}}^{\gamma,\alpha}_{t}$ is adapted to the filtration $\mathbb{F}$. Therefore, there exists a $\mathbb{F}^{\gamma}$-progressively measurable modification of ${\mathcal{L}}^{\gamma,\alpha}$. ∎
For a given $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, the value function of the MFC is defined by,
$v(X_{0},\gamma):=\inf_{\alpha\in{\mathscr{A}}(\gamma)}J^{\gamma}(\alpha,X_{0})=\inf_{\alpha\in{\mathscr{A}}(\gamma)}\mathbb{E}^{\gamma}\left[\int_{0}^{T}F(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},{\mathcal{L}}^{\gamma,\alpha}_{t})\,\mathrm{d}t+G(X^{\alpha}_{\cdot\land T},{\mathcal{L}}^{\gamma,\alpha}_{T})\right]$ (3.1)
where $X^{\alpha}$ solves (SDE$(X_{0},\alpha)$) on $\gamma$. Further, for $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, let $\Gamma(\varrho)$ be the set of all bases $\gamma$ such that there exists $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0}^{\gamma})$ with distribution $\varrho$, let ${\mathcal{I}}^{\gamma}(\varrho)$ be the set of all such initial conditions, and we set
$V(\varrho):=\inf_{\gamma\in\Gamma(\varrho)}\inf_{X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)}v(X_{0},\gamma)=\inf_{\gamma\in\Gamma(\varrho)}\inf_{X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)}\inf_{\alpha\in{\mathscr{A}}(\gamma)}J^{\gamma}(\alpha,X_{0}).$ (3.2)
@@
We say that an MFC problem is invariant under the probability bases, for any $\gamma$, $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$ and $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ we have,
${\mathcal{L}}^{\gamma}(X_{0})=\varrho\quad\Rightarrow\quad v(X_{0},\gamma)=V(\varrho).$
In the diffusion setting, invariance is proved in [23, Theorem 3.6] for ‘rich enough’ probabilistic structures, see [23, Section 2.1.3], and in [29, Proposition 2.4], and our Theorem 7.3 extends it to more general models. Although almost all problems have this property, some natural continuity and growth assumptions are needed as summarized in Assumption 7.2 and the counter example provided in [23, Section 3] highlights their necessity.
@@
3.2 The Potential Mean Field Game of Controls
In this subsection, we give the definition of the potential mean field game of controls equilibrium under common noise; see [15, Definition 2.1] for MFGs under common noise and [26, 27] for general MFC and MFG of controls close to our setting.
In these problems, the representative player takes an $\mathbb{F}$-adapted measure valued process $t\mapsto\nu^{*}_{t}\in{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ as given and solves the optimal control problem $\inf_{\alpha\in{\mathscr{A}}}J_{g}(X_{0},\alpha,\nu)$ where $J_{g}$ is as in (1.4) with cost functions given by (1.3). Then, the Nash equilibrium is defined as follows.
@@ 3.2 The Potential Mean Field Game of Controls
[…]
@@
In the absence of common noise, $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ is a deterministic ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$-valued process. Then, the second condition above reduces to $\nu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})$ for a.e. $t\in[0,T]$. In this case, Borel measurability of $t\mapsto{\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})$ follows directly from the measurability of $(X^{\alpha^{*}},\alpha^{*})$.
@@ 3.3 Main Result
We assume that $F$,$G$ are measurable and for $(t,x,x^{\prime},\nu,a,a^{\prime})\in[0,T]\times{\mathscr{D}}_{n}^{2}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\times U^{2}$ they satisfy:
  • (Differentiability). The functions $F(t,x,a,\cdot)$ and $G(x,\cdot)$ are linearly differentiable with measurable the linear derivatives $\delta_{\nu}F$ and $\delta_{\nu}G$.
  • (Continuity). The maps $F(t,x,a,\cdot)$, $\delta_{\nu}F(t,x,a,\cdot)(x^{\prime},a^{\prime})$, $G(x,\cdot)$ and $\delta_{\nu}G(x,\cdot)(x^{\prime})$ are continuous.
  • (Quadratic Growth).
    $|F(t,x,a,\nu)|+|G(x,\nu)|+$ $\left|\delta_{\nu}F(t,x,a,\nu)(x^{\prime},a^{\prime})\right|+\left|\delta_{\nu}G(x,\nu)(x^{\prime})\right|$
    $\leq C(1+\|x\|_{T}^{2}+|a|^{2}+\mathfrak{s}(\nu)+\|x^{\prime}\|_{T}^{2}+|a^{\prime}|^{2}).$
We prove the main result under the assumption of invariance which is presented in Theorem 7.3 under Assumption 7.2. Additionally, its proof reveals that it holds under a weaker version of invariance as discussed in Remark 4.3 below.